Operations Report – November 25th

Crew 183 Operations Report 25-11-2017

SOL: 3

Name of person filing report: Ashton Stoop

Non-nominal systems:

N/A

Notes on non-nominal systems:

Generator (hours run): 12

Solar— 85%

Diesel – 60%

Propane – 41% volume

Ethanol Free Gasoline (5 Gallon containers for ATV) – 8 gallons remaining

Water (trailer) – 0 gallons added to static

Water (static) – 325 gallons

Trailer to Static Pump used – No

Water (loft) – Static to Loft Pump used – yes

Water Meter: 48

Toilet tank emptied: No

ATV’s Used: Honda, 350.1

Oil Added? No

ATV Fuel Used: 2 Gals

# Hours the ATVs were Used today: 1

Notes on ATVs: Everything ran smoothly today.

Deimos rover used: Yes

Hours: 1

Beginning charge: 100%

Ending charge: 95%

Currently charging: Yes

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: NO

Hours:

Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used: NO

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used: NO

Hours:

Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where? NO

General notes and comments:

It’s hot in those EVA suits…

Summary of internet:

Every evening after dinner time it slows down a lot.

Summary of suits and radios:

There have been some issues with radios and headsets over the past few days. We will need to do a test and see which ones are working properly and which aren’t. Some of the EVA suits are having issues. Suit 4 only had one fan working and suit 1 wasn’t working at all. We will double check all the charges to see if that was an issue. One of the XLR cables isn’t working properly. We’re going to measure the current going through and give a more detailed diagnosis in the near future.

Summary of Hab operations:

Things are operating smoothly.

Summary of GreenHab operations:

Everything is growing well and grow lights are operational.

Summary of ScienceDome operations:

Sedementation tests are proceeding as planned. We received a new soil solidity device today.

Summary of RAMM operations:

Not used. It is not yet fully operational.

Summary of health and safety issues:

Waiting and Carbon monoxide detectors.

Questions, concerns and requests to Mission Support:

-Things are going well overall.